PR 9-1A Entries related to uncollectible accounts year ended December 31 Jan. 29, Received 35% of the $9,000 balance owed by Kovar Co., a bankrupt business, Apr. 18. Reinstated the account of Spencer Clark, which had been written off in the OBJ.4 3.$1,390,000 The following transactions were completed by Daws Company during the current fiscal General Ledger and wrote off the remainder as uncollectible Show Me How preceding year as uncollectible. Journalized the receipt of $4,000 cash in full payment of Clarks account. Aug. 9. Wrote off the $11,850 balance owed by Iron Horse Co., which has no assets. Nov. 7. Reinstated the account of Vinyl Co., which had been written off in the preceding year as uncollectible. Journalized the receipt of $7,000 cash in full payment of the account. $12,100; DeVine Co., $8,110; Moser Distributors, $21,950; Oceanic Optics, $10,000. that $60,000 will be uncollectible. Journalized the adjusting entry Dec. 31. Wrote off the following accounts as uncollectible (one entry): Beth Connelly Inc. 31. Based on an analysis of the $1,450,000 of accounts receivable, it was estimated Instructions 1. Record the January 1 credit balance of $54,200 in a T account for Allowance for Doubtful Accounts. Journalize the transactions. Post each entry that affects the following selected T accounts and determine the new balances: 2. Allowance for Doubtful Accounts Bad Debt Expense 3. Determine the expected net realizable value of the accounts receivable as of December 31. 4. Assuming that instead of basing the provision for uncollectible accounts on an analysis of receivables the adjusting entry on December 31 had been based on an estimated expense oftá of 1% of the sales of $13,200,000 for the year, determine the following: a. Bad debt expense for the year b. Balance in the allowance account after the adjustment of December 31 c. Expected net realizable value of the accounts receivable as of December 31
PR 9-4A Details of notes receivable and related entries Flush Mate Co. wholesales bathroom fixtures. During the current fiscal year, Flush Mate Co. received the following notes: OBJ. 6 1. Note 2: Due date, June 22; Interest due at maturity, $360 Date Mar. 6 Apr. 23 July 20 Sept. 6 Nov. 29 Dec. 30 Face Amount Interest Rate Term 80,000 24,000 42,000 54,000 27,000 72,000 45 days 60 days 120 days 90 days 60 days 30 days 2. 3. How 5. 6. Instructions 1. Determine for each note (a) the due date and (b) the amount of interest due at maturity, identifying each note by number 2. Journalize the entry to record the dishonor of Note (3) on its due date 3. Journalize the adjusting entry to record the accrued interest on Notes (5) and (6) on 4. Journalize the entries to record the receipt of the amounts due on Notes (5) and (6) in January
Answer #1

Since, multiple questions have been posted, I have answered all the parts of PR 9-1A.

_______

We will have to first prepare the journal entries (which is Part 2) and thereafter Complete T-Accounts as required in Part 1 and Part 2.

_____

Part 2)

The journal entries are prepared as below:

Date Account Name Debit Credit
Jan.29 Cash (9,000*35%) $3,150
Allowance for Doubtful Accounts (9,000*65%) $5,850
Accounts Receivable-Kovar Co. $9,000
Apr.18 Accounts Receivable-Spencer Clark $4,000
Allowance for Doubtful Accounts $4,000
Apr.18 Cash $4,000
Accounts Receivable-Spencer Clark $4,000
Aug.9 Allowance for Doubtful Accounts $11,850
Accounts Receivable-Iron Horse Co. $11,850
Nov.7 Accounts Receivable-Vinyl Co. $7,000
Allowance for Doubtful Accounts $7,000
Nov.7 Cash $7,000
Accounts Receivable-Vinyl Co. $7,000
Dec.31 Allowance for Doubtful Accounts $52,160
Accounts Receivable-Beth Connelly Inc. $12,100
Accounts Receivable-DeVine Co. $8,110
Accounts Receivable-Moser Distributors $21,950
Accounts Receivable-Oceanic Optics $10,000
Dec.31 Bad Debt Expense (60,000 + 4,660) $64,660
Allowance for Doubtful Accounts $64,660

_____

Part 1 and 2)

The T-Accounts are given as follows:

Allowance for Doubtful Accounts (T-Account)
Jan.29 5,850 54,200 Jan.1
Aug.9 11,850 4,000 Apr.18
Dec.31 52,160 7,000 Nov.7
Dec.31 (Unadjusted Balance) 4,660 64,660 Dec.31 (Adjusting Entry)
$60,000 Dec.31 (Adjusted Bal.)

_____

Bad Debt Expense
Dec.31 (Adjusting Entry) 64,660

_____

Part 3)

The expected net realizable value of the accounts receivable as of December 31 is determined as below:

Expected Net Realizable Value = Total Value of Accounts Receivable - Estimated Uncollectible = 1,450,000 - 60,000 = $1,390,000

_____

Part 4)

a)

The value of bad debt expense for the year is calculated as follows:

Bad Debt Expense = Net Sales*1%*1/2 = 13,200,000*1%*1/2 = $66,000

_____

b)

The balance in the allowance account after the adjustment of December 31 is arrived as below:

Balance in Allowance Account = Bad Debt Expense - Unadjusted Balance = 66,000 - 4,660 = $61,340

_____

c)

The expected net realizable value of the accounts receivable as of December 31 is determined as follows:

Expected Net Realizable Value = Total Value of Accounts Receivable - Balance in Allowance Account = 1,450,000 - 61,340 = $1,388,660
Answer #2
PR9-1A Entries related to uncollectible accounts OBJ. 4 The following transactions were completed by Emmanuel Company during the current fiscal year ended December 31: → SHOW ME HOW Jan. 29. Received 40% of the $17,000 balance owed by Jankovich Co., a bankrupt busi- ness, and wrote off the remainder as uncollectible. Apr. 18. Reinstated the account of Vince Karm, which had been written off in the preced- ing year as uncollectible. Journalized the receipt of $7,560 cash in full payment of Karm's account. Aug. 9. Wrote off the $22,380 balance owed by Golden Stallion Co., which has no assets. Nov. 7. Reinstated the account of Wiley Co., which had been written off in the preced- ing year as uncollectible. Journalized the receipt of $13,220 cash in full payment of the account. Dec. 31. Wrote off the following accounts as uncollectible (one entry): Claire Moon Inc., $22,860; Jet Set Co., $15,320; Randall Distributors, $41,460; Harmonic Audio, $18,890. 31. Based on an analysis of the $2,740,000 of accounts receivable, it was estimated that $113,330 will be uncollectible. Journalized the adjusting entry. Instructions 1. Record the January 1 credit balance of $102,380 in a T account for Allowance for Doubtful Accounts. 2. Journalize the transactions. Post each entry that affects the following selected T accounts and determine the new balances: Allowance for Doubtful Accounts Bad Debt Expense 3. Determine the expected net realizable value of the accounts receivable as of December 31. 4. Assuming that instead of basing the provision for uncollectible accounts on an analysis of receivables the adjusting entry on December 31 had been based on an estimated expense of 12 of 1% of the sales of $24,900,000 for the year, determine the following: a. Bad debt expense for the year. b. Balance in the allowance account after the adjustment of December 31. c. Expected net realizable value of the accounts receivable as of December 31.